OpenStudy (anonymous):

The Everton College store paid $1794 for an order of 46 calculators. The store paid $10 for each scientific calculator. The others, all graphing calculators, cost the store $56 each. How many of each type of calculator was ordered?

OpenStudy (anonymous):

let x= $10, y=$ 56 x+y=46 y= 46-x 10x+56y=1794 substitution y= 46-x and solve

OpenStudy (anonymous):

help

OpenStudy (anonymous):

10x+56(46-x)=1794 can u solve from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

10x+2.576-56x=1794 -46x =1794-2.576 x=-782/-46 x=17

OpenStudy (anonymous):

i need to figure how many of each type was ordered

OpenStudy (anonymous):

substitution x= 17 and solve for y

OpenStudy (anonymous):

x+y=46 17+y=46 y = 46-17 y = 29 $56 has 29 calculators. $10 has 17 calculators.

OpenStudy (anonymous):

tty